# Weaving the (AdS) spaces with partial entanglement entropy threads

### v3

@@ Abstract
In the context of the AdS/CFT correspondence, we propose a general scheme for reconstructing bulk geometric quantities in a static pure AdS background using the partial entanglement entropy (PEE), a measure of the entanglement structure on the boundary CFT. The PEE between any two points $\mathcal{I}(\textbf{x},\textbf{y})$ serves as the fundamental building block of the PEE structure. Following [1], we geometrize any two-point PEE $\mathcal{I}(\textbf{x},\textbf{y})$ by the bulk geodesic connecting two boundary points x and y, which we call the PEE thread, with the density of the threads determined by the boundary PEE structure. In the AdS bulk, the set of all the PEE threads forms a continuous “network”, which we call the PEE network.
In this paper, we show that the density of the PEE threads passing through any bulk point is exactly $1/(4G)$. Based on this observation we give a complete reformulation of the Ryu-Takayanagi (RT) formula for a generic boundary region in general dimensional Poincaré AdS space. More explicitly, for any static boundary region $A$, the homologous surface $\Sigma_{A}$ that has the minimal number of intersections with the bulk PEE network is exactly the RT surface of $A$, and the minimal number of intersections reproduces the holographic entanglement entropy. The reconstruction for the area of bulk geometric quantities by counting the number of intersections with the bulk PEE network applies to generic bulk geometric quantities. Interestingly, this reconstruction indicates a pure geometric statement, which is exactly the so-called Crofton formula in Poincaré AdS.
@@ 1 Introduction
The AdS/CFT correspondence [2, 3, 4] states that the quantum theory of gravity in asymptotic AdSd+1 spacetime is equivalent to a certain CFTd on the asymptotic boundary. This provides a window into understanding both the classical and quantum aspects of gravitational theories based on the information in the boundary CFT, using the dictionary of the correspondence. Among the studies of the correspondence, holographic entanglement entropy plays a central role, as it relates boundary quantum entanglement structure to the geometry of spacetime. Several important developments have been made along this line [5, 6, 7, 8, 9, 10, 11, 12, 13]. These achievements began with the Ryu-Takayanagi (RT) formula [5] which links the entanglement entropy of any boundary region and the area of the bulk minimal surface homologous to that boundary region. This proposal was later refined into the covariant version [6, 13] and the version including quantum corrections [8, 9, 11, 12]. For more recent developments on the bulk reconstruction inspired by holographic study of quantum entanglement, one consults the following review papers [14, 15].
The possibility to reconstruct the bulk geometry from the entanglement structure of the boundary field theory was soon realized after the RT formula was proposed, see [16, 17] for the earliest discussions. In this paper, we will focus on the explicit reconstruction for the area of bulk geometric quantities in terms of boundary entanglement structure measures. So far, several approaches have been explored for this goal. For example, the reconstruction of certain bulk curves via the differential entropy [18, 19, 20, 21, 22, 23] by studying the geodesics tangent to the curve, the reformulation of the RT formula as the maximal flux of the bit threads in AdS space out of the region [24, 25, 26], and the simulation of the AdS space based on the tensor networks [27, 28, 29, 30, 31, 32, 33, 34, 35] where the RT surface is interpreted as the homologous path in the network with the minimal number of cuts. See [36] for a detailed review on the above approaches.
@@
In section 2, we will briefly review the PEE in the vacuum state of a holographic CFT, and the setup [1] for its geometrization as the PEE threads in the AdS bulk. Also we will go through the calculations [1] that count the number of intersections between the RT surfaces of spherical regions and the bulk PEE network, which is very useful for our later discussions. In section 3, we compute the number of intersections between the RT surfaces for generic boundary regions and the bulk PEE network. We show that for any given boundary region, the RT surface is reproduced by identifying the homologous surface that has the minimal number of intersections with the bulk PEE network. Furthermore, this minimal number of intersections is exactly the holographic entanglement entropy given by the RT formula. This gives a complete reformulation of the RT formula, which is quite similar to the calculation of the entanglement entropy in tensor network models. In section 4, we use the PEE threads to reconstruct a generic bulk co-dimension two surface by counting the number of intersections between the surface and the PEE network. In section 5, we show that our reconstruction scheme is equivalent to a pure mathematical statement, which is precisely the well-known Crofton formula. In the last section, we give a summary on our results and give some possible interesting future directions.
@@ 2.1 Partial entanglement entropy
As was shown in [61, 40], the above requirements have a unique solution for states with Poncaré symmetry. Furthermore, for the vacuum state of a CFT on a plane, the formula of the solution is totally determined by the above requirements.
According to the properties of additivity and permutation symmetry, the PEE structure are fully described by two-point PEEs $\mathcal{I}(\textbf{x},\textbf{y})$, and $\mathcal{I}(A,B)$ can be written as a double integral over $A$ and $B$
@@
where $\Omega_{d-2}={2\pi^{\frac{d-1}{2}}}/{\Gamma\left(\frac{d-1}{2}\right)}$ is the area of $(d-2)$-dimensional unit sphere. One can either derive the above formula for two-point PEE via the solution to all the physical requirements [61, 40], or using the so-called additive-linear-combination (ALC) proposal in quasi-one-dimensional system to construct PEE [42]. See [1] for the details about the derivation of (2). Although the PEE structure (2) may not capture all the information of the entanglement structure in a CFT, we will see that it is enough to reconstruct the geometric quantities at order $\mathcal{O}(c)$ in the gravity side of AdS/CFT.
@@
where $A\cup\bar{A}$ makes up a pure state and $\epsilon$ represents a regularization cutoff. Note that, the requirement of normalization is quite subtle as it is an equality between two divergent quantities which are normalized in different schemes, see [42] for more discussions on this requirement. We should keep in mind that, we only impose the normalization requirement to spherical regions, where the relation between the geometrical cutoff and the UV cutoff can be explored, to get the solution (2). The solution may not exist if we impose the normalization requirement to generic regions (especially disconnected regions) [^1] Nevertheless, in CFTd entanglement entropies for various shapes of connected regions have been carried out based on (2) and (3) [62, 63, 64, 65], which are in good (but not exact) agreement with the results derived from other methods. In these papers, the authors studied the mutual information that satisfies additivity (EMI), which coincides with the PEE [40] in these scenarios. See [40, 42] for discussion on the relationship between the PEE and the EMI.. As implied in [1], giving up this requirement for non-spherical regions is the key for our scheme to reconstruct the geometric quantities in AdS.
@@ 2.2 PEE threads and the partial entanglement network
[…]
  • the number of PEE threads connecting any two boundary non-overlapping regions (or boundary points) $A$ and $B$ should match the PEE $\mathcal{I}(A,B)$.
@@
Figure 1: Visualizations of PEE threads (gray) on a time slice of Poincaré AdS4 and global AdS3, where the boundary sites (on the plane) are discretized. Left: the PEE threads with larger radii are denoted with thinner curves, whose thickness reflects the density of the PEE threads. Right: with appropriate discretizations, they form a hyperbolic tiling.
@@
$\mathrm{d}s^{2}=\frac{-\mathrm{d}t^{2}+\mathrm{d}z^{2}+\mathrm{d}r^{2}+r^{2}\mathrm{d}\Omega_{d-2}^{2}}{z^{2}}\,,$ (4)
where $\mathrm{d}\Omega_{d-2}^{2}=\mathrm{d}\phi_{1}^{2}+\sin^{2}\phi_{1}\mathrm{d}\phi_{2}^{2}+\cdots+\sin^{2}\phi_{1}\cdots\sin^{2}\phi_{d-3}\mathrm{d}\phi_{d-2}^{2}.$ In higher dimensions, due to the rotational symmetry, we restrict to a 2-dimensional slice with $\phi_{i}=0$. Since the PEE threads are just the bulk geodesics anchored on $O$, the vector field $V_{O}^{\mu}(Q)$ is tangent to these geodesics, such that
@@
is the unit vector tangent to the geodesics emanating from $O$, and $z,~{}r$ are the coordinates of the bulk point $Q$ on the $\phi_{i}=0$ slice. The norm $|V_{O}(Q)|$ is then settled down by the requirement that, the number of PEE threads from $O$ to any boundary region $\mathrm{d}y$ should match the PEE $\mathcal{I}(0,\mathrm{d}y)$, where $y$ is the radial coordinate on the boundary.
$O$$\mathrm{d}y$$\mathrm{d}\Sigma$$-\ell$$\ell$$y$$Q$$\theta$
@@
$\mathcal{F}(V_{O}^{\mu},\mathrm{d}\Sigma)=$ $\mathrm{d}\theta\mathrm{d}\Omega_{d-2}\sqrt{h}V_{O}^{\mu}(\theta)n_{\Sigma,\mu}(\theta)$ (7)
$=$ $\mathrm{d}\theta\mathrm{d}\Omega_{d-2}\frac{\bar{r}^{d-2}}{\bar{z}^{d-2}}\sqrt{h_{\theta\theta}}|V_{O}(\theta)|\sin\theta$
$=$ $\mathrm{d}\theta\mathrm{d}\Omega_{d-2}\sqrt{h_{\theta\theta}}|V_{O}(\theta)|\cot^{d-2}\theta\sin\theta,$
@@
  • Theorem 1: the density of intersections between the bulk PEE network and the RT surface $\mathcal{E}_{A}$ for any spherical region (or single interval) $A$ is given by the constant $\frac{1}{4G}$ everywhere on $\mathcal{E}_{A}$.
Later, we denote the number of intersections between any bulk co-dimension-two surface $\Sigma$ and the PEE network as $\mathcal{N}(\Sigma)$. So we have that, for spherical regions $A$
@@ 2.3 The weight of the PEE threads
$S_{A}=\int_{\Sigma_{A}}\mathrm{d}\Sigma_{A}\sqrt{h}V_{A}^{\mu}n_{\mu}.$ (16)
Naively, we expect that (16) should reproduce the same entanglement entropy as the RT formula for an arbitrary region $A$. Nevertheless, this is not true for non-spherical regions, see [1] for explicit analysis when $A$ is a strip region. To fix this problem, we should modify (3) in some way. The discussion for the two-interval case in [1] in CFT2 provides a clear clue for the modification.
@@
Figure 3: This figure is extracted from [1]. Here $A=A_{1}\cup A_{2}$, $\bar{A}=\bar{A}_{1}\cup\bar{A}_{2}$ and $A\bar{A}$ is in the vacuum state of the holographic CFT2. Representative PEE threads (dashed curves) in Poincaré AdS3. Each PEE thread crosses the disconnected RT surface for different number of times, which represents different weight.
Let us consider a two interval region $A=A_{1}\cup A_{2}$ whose RT surface is given by the blue curves, see Fig. 3. If we naively apply (3), we should only count the threads connecting $A_{i}$ and $\bar{A}_{j}$, which gives $S_{A}=\mathcal{I}(A_{1},\bar{A})+\mathcal{I}(A_{2},\bar{A})$. However, the RT formula implies
@@
The above observation is consistent with (3) for single interval cases where only the threads connecting the interval $A$ and its complement $\bar{A}$ have non-zero weight $\omega(A,\bar{A})=1$. From bulk PEE network perspective, it is important to note that, (19) indeed counts $\mathcal{N}(\mathcal{E}_{A})$, the number of intersections between the RT surface $\mathcal{E}_{A}$ and the bulk PEE network.
@@ 3 Reformulation of the RT formula
One of the main task of this paper and [1] is to give a complete reformulation for the RT formula based on the PEE threads setup. As we have just reviewed, for static spherical regions and multi-intervals in the vacuum of holographic CFT, the area of the RT surfaces can be reproduced by counting the number of intersections $\mathcal{N}(\mathcal{E}_{A})$. Nevertheless, the discussion in [1] only reproduces the holographic entanglement entropy for spherical regions (plus multi-intervals in 2-dimensions) provided the RT surface $\mathcal{E}_{A}$ is known to us.
To fulfill our task, our strategy is to prove the following statement, which is the key to give our complete reformulation of the RT formula.
  • Theorem 2: given a generic boundary region $A$ and an arbitrary surface $\Sigma_{A}$ homologous to $A$, the density of intersections between $\Sigma_{A}$ and the bulk PEE network everywhere on $\Sigma_{A}$ is also given by $\frac{1}{4G}$.
@@
And the RT surface $\mathcal{E}_{A}$ can be identified by finding the homologous surface that minimizes the number of intersections $\mathcal{N}(\Sigma_{A})$ among all possible $\Sigma_{A}$. Finally, provided the theorem 2 the RT formula is completely reformulated by
@@
In other words, given an arbitrary boundary region $A$, the RT surface is the $\Sigma_{A}$ that has the minimal number of intersections with the PEE network. This number is exactly the holographic entanglement entropy $S_{A}$ calculated by the RT formula. The remaining task in this section is to prove the theorem 2.
[…]
$\mathcal{N}(\Sigma_{A})=\frac{1}{2}\int_{\partial\mathcal{M}}\mathrm{d}^{d-1}\textbf{x}\int_{\partial\mathcal{M}}\mathrm{d}^{d-1}\textbf{y}~{}\omega_{\Sigma_{A}}(\textbf{x},\textbf{y})\mathcal{I}(\textbf{x},\textbf{y}),$ (22)
where the integration domain of x and y is the whole AdS boundary $\partial\mathcal{M}=A\bar{A}$. Here we have considered all the PEE threads and their weights instead of only those connecting $A$ and $\bar{A}$. The factor $1/2$ appears as we count both $\mathcal{I}(\textbf{x},\textbf{y})$ and $\mathcal{I}(\textbf{y},\textbf{x})$ in the integration, hence every PEE thread has been counted twice. Conducting the integration (22) is formidable since it is impossible to count the weight $\omega_{\Sigma_{A}}(\textbf{x},\textbf{y})$ of all the PEE threads and express it using a closed function. The formula (22) will be quite useful when we relate our results to the Crofton formula.

### v4

# Weaving the AdS spaces with partial-entanglement-entropy threads

@@ Abstract
In the context of the AdS/CFT correspondence, we propose a general scheme for reconstructing bulk geometric quantities in a static pure AdS background using the partial entanglement entropy (PEE), a measure of the entanglement structure on the boundary CFT. The PEE between any two points $\mathcal{I}(\textbf{x},\textbf{y})$ serves as the fundamental building block of the PEE structure. Any two-point PEE $\mathcal{I}(\textbf{x},\textbf{y})$ can be geometrized by the bulk geodesic connecting two boundary points x and y, which we call the PEE thread, with the density of the threads determined by the boundary PEE structure. In the AdS bulk, the set of all the PEE threads forms a continuous “network”, which we call the PEE network.
In this paper, we show that the density of the PEE threads passing through any bulk point is exactly $1/(4G)$. Based on this observation we give a reformulation of the Ryu-Takayanagi (RT) formula for a generic boundary region in general dimensional Poincaré AdS space. More explicitly, for any static boundary region $A$, the homologous surface $\Sigma_{A}$ that has the minimal number of intersections with the bulk PEE network is exactly the RT surface of $A$, and the minimal number of intersections reproduces the holographic entanglement entropy. The reconstruction for the area of bulk geometric quantities by counting the number of intersections with the bulk PEE network applies to generic bulk geometric quantities. Interestingly, this reconstruction indicates a pure geometric statement, which is exactly the so-called Crofton formula in Poincaré AdS.
keywords:
AdS-CFT Correspondence, Gauge-Gravity Correspondence, Models of Quantum Gravity
articletype: Paper
@@ 1 Introduction
The AdS/CFT correspondence [73, 42, 94] states that the quantum theory of gravity in asymptotic AdSd+1 spacetime is equivalent to a certain CFTd on the asymptotic boundary. This provides a window into understanding both the classical and quantum aspects of gravitational theories based on the information in the boundary CFT, using the dictionary of the correspondence. Among the studies of the correspondence, holographic entanglement entropy plays a central role, as it relates boundary quantum entanglement structure to the geometry of spacetime. Several important developments have been made along this line [80, 60, 64, 39, 37, 35, 33, 32]. These achievements began with the Ryu-Takayanagi (RT) formula [80] which links the entanglement entropy of any boundary region and the area of the bulk minimal surface homologous to that boundary region. This proposal was later refined into the covariant version [60, 32] and the version including quantum corrections [64, 39, 35, 33]. For more recent developments on the bulk reconstruction inspired by holographic study of quantum entanglement, one consults the following review papers [15, 38].
The possibility to reconstruct the bulk geometry from the entanglement structure of the boundary field theory was soon realized after the RT formula was proposed, see [84, 85] for the earliest discussions. In this paper, we will focus on the explicit reconstruction for the area of bulk geometric quantities in terms of boundary entanglement structure measures. So far, several approaches have been explored for this goal. For example, the reconstruction of certain bulk curves via the differential entropy [6, 53, 25, 30, 27, 28] by studying the geodesics tangent to the curve, the reformulation of the RT formula as the maximal flux of the bit threads in AdS space out of the region [40, 50, 52, 51, 47, 48, 3, 4, 79, 68, 34], and the simulation of the AdS space based on the tensor networks [83, 36, 43, 82, 77, 76, 49, 13, 14] where the RT surface is interpreted as the homologous path in the network with the minimal number of cuts. See [23] for a detailed review on the above approaches.
@@
This paper is based on the geometric scheme in [66], where the geometric counterpart of PEE, namely PEE threads which are geodesics in the time slice, is introduced. Bit-thread configurations and entanglement entropies for spherical symmetrical regions are recovered by PEE threads. A key observation in [66] from AdS3 is that the contribution of a PEE to holographic entanglement entropy is weighted by the number of intersections of its PEE thread with the RT surface. In this paper, we provide a proof of this statement in Poincaré AdS, and reformulate the area of RT surface, i.e. holographic entanglement entropy, as a double integral. And we show that our reformulation is nothing but the Crofton formula, and shall be able to extend to general spacetimes.
In section 2, we will briefly review the PEE in the vacuum state of a holographic CFT, and the setup [66] for its geometrization as the PEE threads in the AdS bulk. Also we will go through the calculations [66] that count the number of intersections between the RT surfaces of spherical regions and the bulk PEE network, which is very useful for our later discussions. In section 3, we compute the number of intersections between the RT surfaces for generic boundary regions and the bulk PEE network. We show that for any given boundary region, the RT surface is reproduced by identifying the homologous surface that has the minimal number of intersections with the bulk PEE network. Furthermore, this minimal number of intersections is exactly the holographic entanglement entropy given by the RT formula. This gives a reformulation of the RT formula, which is quite similar to the calculation of the entanglement entropy in tensor network models. In section 4, we use the PEE threads to reconstruct a generic bulk co-dimension two surface by counting the number of intersections between the surface and the PEE network. In section 5, we show that our reconstruction scheme is equivalent to a pure mathematical statement, which is precisely the well-known Crofton formula. In the last section, we give a summary on our results and give some possible interesting future directions.
@@ 2 The partial entanglement network
In this section, we first introduce the basic aspects of PEE. Then we review the main results of [66], which geometrizes the two-point PEE structure into bulk geodesics (namely the PEE threads), and propose to weight the PEE threads differently when conputing holographic entanglement entroy. These results set the stage for our next section, where a reformulation of the RT formula in Poincaré AdS will be provided.
@@ 2.1 Partial entanglement entropy
As was shown in [92, 21], the above requirements have a unique solution for states with Poincaré symmetry. Furthermore, for the vacuum state of a CFT on a plane, the formula of the solution is totally determined by the above requirements.
According to the properties of additivity and permutation symmetry, the PEE structure is fully described by two-point PEEs $\mathcal{I}(\textbf{x},\textbf{y})$, and $\mathcal{I}(A,B)$ can be written as a double integral over $A$ and $B$
@@
where $\Omega_{d-2}={2\pi^{\frac{d-1}{2}}}/{\Gamma\left(\frac{d-1}{2}\right)}$ is the area of (d2)(d-2)-dimensional unit sphere. One can either derive the above formula for two-point PEE via the solution to all the physical requirements [21, 92], or use the so-called additive-linear-combination (ALC) proposal in quasi-one-dimensional system to construct PEE [45]. See [66] for the details about the derivation of (2). Although the PEE structure (2) may not capture all the information of the entanglement structure in a CFT, we will see that it is enough to reconstruct the geometric quantities at order $\mathcal{O}(c)$ in the gravity side of AdS/CFT.
@@
where $A\cup\bar{A}$ makes up a pure state and $\epsilon$ represents a regularization cutoff. Note that, the requirement of normalization is quite subtle as it is an equality between two divergent quantities which are normalized in different schemes, see [45] for more discussions on this requirement. We should keep in mind that, we only impose the normalization requirement to spherical regions for the following reasons. Firstly, imposing this reqruirement to spherical regions is enough to determine the two-point PEE structure (2). Secondly, if we futher impose this requirement to more generic regions, the solution does not exist. In other words, the entanglement entropy for non-spherical regions can not be reproduced by (2) and (3)  [^1] In [19, 18, 17, 16], the entanglement entropies for various shapes of connected regions for the vacuum state of several CFTs have been carried out based on (2) and (3). Compared with the results derived from the RT formula, these results have a similar formula, but the coefficents at each order are different. In these papers, the authors studied the mutual information that satisfies additivity (EMI), which coincides with the PEE [92] in these scenarios. See [92, 45] for discussion on the relationship between the PEE and the EMI.. One can understand this more clearly in the case of one dimensional lattice circle, where the spherical regions are single intervals. It was pointed out in [92] that, the number of single intervals equals to the number of two-point PEEs, hence the number of equations (3) equals the number of two-points PEEs. Thirdly, as implied in [66], if we want to reproduce the RT formula for multi-intervals, we need to give up this requirement. Later we will see that, the holographic entanglement $S_{A}$ for non-spherical regions can be reproduced from the two-point PEE structure by including contributions from other two-point PEEs, except those represented by $\mathcal{I}(A,\bar{A})$.
@@ 2.2 PEE threads and the partial entanglement network
In AdS/CFT, we introduced a scheme to geometrize the PEEs in [66], where the boundary two-point PEEs $\mathcal{I}(\textbf{x},\textbf{y})$ are represented by the bulk geodesics connecting two boundary points $\{\textbf{x},\textbf{y}\}$, which we call the PEE threads [66]. This geometrization looks similar to the bit thread configurations. Nevertheless they are quite different objects, for example the PEE thread configuration is totally determined by the boundary state and they intersect with each other, while the bit thread configuration are highly degenerate and bit threads do not intersect with each other. We only consider the vacuum state of the CFTd and a static time slice in AdSd+1. The PEE threads emanating from any point x can be represented by a divergenceless vector field $V_{\textbf{x}}^{\mu}=|V_{\textbf{x}}|\tau^{\mu}$, where $\tau^{\mu}$ is the unit vector tangent to the geodesics emanating from x. The norm $|V_{\textbf{x}}|$ characterizes the density of the threads, which is determined by the requirement that,
  • setup: the number of PEE threads connecting any two boundary non-overlapping regions (or boundary points) $A$ and $B$ should match the PEE $\mathcal{I}(A,B)$.
@@
Figure 1: Visualizations of PEE threads (gray) on a time slice of Poincaré AdS4 and global AdS3, where the boundary sites (on the plane) are discretized. Upper: the PEE threads with larger radii are denoted with thinner curves, whose thickness reflects the density of the PEE threads. Bottom: with appropriate discretizations, they form a hyperbolic tiling.
@@
$\mathrm{d}s^{2}=\frac{-\mathrm{d}t^{2}+\mathrm{d}z^{2}+\mathrm{d}r^{2}+r^{2}\mathrm{d}\Omega_{d-2}^{2}}{z^{2}}\,.$ (4)
In higher dimensions, due to the rotational symmetry, we restrict to a 2-dimensional slice with $\phi_{i}=0$. Since the PEE threads are just the bulk geodesics anchored on $O$, the vector field $V_{O}^{\mu}(Q)$ is tangent to these geodesics, such that
@@
is the unit vector tangent to the geodesics emanating from $O$, and z,rz,~r are the coordinates of the bulk point $Q$ on the $\phi_{i}=0$ slice. The norm $|V_{O}(Q)|$ is then settled down by the requirement that, the number of PEE threads from $O$ to any boundary region $\mathrm{d}y$ should match the PEE $\mathcal{I}(0,\mathrm{d}y)$, where $y$ is the radial coordinate on the boundary.
[…]
@@
$\mathcal{F}(V_{O}^{\mu},\mathrm{d}\Sigma)=\mathrm{d}\theta\mathrm{d}\Omega_{d-2}\sqrt{h_{\theta\theta}}|V_{O}(\theta)|\cot^{d-2}\theta\sin\theta,$ (7)
@@
  • Statement 1: the density of intersections between the bulk PEE network and the RT surface $\mathcal{E}_{A}$ for any spherical region (or single interval) $A$ is given by the constant $\frac{1}{4G}$ everywhere on $\mathcal{E}_{A}$.
Later, we denote the number of intersections between any bulk co-dimension-two surface $\Sigma$ and the PEE network as $\mathcal{N}(\Sigma)$. So we have that, for spherical regions AA
@@ 2.3 The weight of the PEE threads
${\mathcal{I}(A,\bar{A})}=\int_{\Sigma_{A}}\mathrm{d}\Sigma_{A}\sqrt{h}V_{A}^{\mu}n_{\mu}.$ (16)
Naively, we expect that (16) should reproduce the same entanglement entropy as the RT formula for an arbitrary region $A$. Nevertheless, this is not true for non-spherical regions, see [66] for explicit analysis when $A$ is a strip region. This implies that the normalization property only applies to spherical regions in Poincaré AdS, and (3) should be modified to reproduce the holographic entanglement entropy for more generic boundary regions.
$\bar{A}_{1}$$A_{2}$$A_{1}$$\bar{A}_{2}$$\bar{A}_{2}$$\omega(A_{1},A_{2})=2$$\omega(\bar{A}_{1},\bar{A}_{2})=0$$\omega(A,\bar{A})=1$
@@
The discussion for the two-interval case in [66] in CFT2 provides a clear clue for the modification. Let us consider a two interval region $A=A_{1}\cup A_{2}$ whose RT surface is given by the blue curves, see Fig. 3. If we naively apply (3), we should only count the threads connecting $A_{i}$ and $\bar{A}_{j}$, which gives $S_{A}=\mathcal{I}(A_{1},\bar{A})+\mathcal{I}(A_{2},\bar{A})$. However, the RT formula implies
@@
From bulk PEE network perspective, it is important to note that, (19) indeed counts $\mathcal{N}(\mathcal{E}_{A})$, the number of intersections between the RT surface $\mathcal{E}_{A}$ and the bulk PEE network.
@@ 3 Reformulation of the RT formula
One of the main task of this paper is to give a reformulation for the RT formula based on the PEE threads setup. As we have just reviewed [66], the area of the RT surfaces can be reproduced by counting the number of intersections $\mathcal{N}(\mathcal{E}_{A})$ between the RT surface $\mathcal{E}_{A}$ and the PEE network. Nevertheless, this observation is still far from a reformulation of the RT formula for the following reasons.
  • Firstly, the discussions in [66] were based on the assumption that we already know the RT surface, and no method was provided to identify the RT surface.
  • Secondly, the computation for $\mathcal{N}(\mathcal{E}_{A})$ in [66] was performed only for the RT surfaces of spherical regions in general dimensions, and multi-intervals in 2-dimensions. In order to give a reformulation for the RT formula, we need to compute the $\mathcal{N}(\mathcal{E}_{A})$ for generic RT surfaces.
To fulfill our task, our strategy is to prove the following statement, which is the key to give our reformulation of the RT formula.
  • Statement 2: in Poincaré AdS space, given a generic boundary region $A$ and an arbitrary surface $\Sigma_{A}$ homologous to $A$, the density of intersections between $\Sigma_{A}$ and the bulk PEE network everywhere on $\Sigma_{A}$ is also given by $\frac{1}{4G}$.
@@
And the RT surface $\mathcal{E}_{A}$ can be identified by finding the homologous surface that minimizes the number of intersections $\mathcal{N}(\Sigma_{A})$ among all possible $\Sigma_{A}$. Finally, provided the statement 2 the RT formula is reformulated by
@@
In other words, given an arbitrary boundary region $A$, the RT surface is the $\Sigma_{A}$ that has the minimal number of intersections with the PEE network. This number is exactly the holographic entanglement entropy $S_{A}$ calculated by the RT formula. The remaining task in this section is to prove the statement 2.
Before we prove the statement 2, we give two explicit formulas to calculate the number of intersections $\mathcal{N}(\Sigma_{A})$ for a generic homologous surface. In a time slice of Poincaré AdSd+1 in general dimensions, let us consider an arbitrary boundary region $A$ (connected or disconnected) and an arbitrary co-dimension two surface $\Sigma_{A}$ homologous to $A$. Similarly we define the weight of a PEE thread as the number of times it intersects with the surface $\Sigma_{A}$, and write it as $\omega_{\Sigma_{A}}(\textbf{x},\textbf{y})$, where $\{\textbf{x},\textbf{y}\}$ are the boundary points where the PEE threads anchor. Then the number of intersections $\mathcal{N}(\Sigma_{A})$ is calculated by
$\boxed{\mathcal{N}(\Sigma_{A})=\frac{1}{2}\int_{\partial\mathcal{M}}\mathrm{d}^{d-1}\textbf{x}\int_{\partial\mathcal{M}}\mathrm{d}^{d-1}\textbf{y}~\omega_{\Sigma_{A}}(\textbf{x},\textbf{y})\mathcal{I}(\textbf{x},\textbf{y}),}$ (22)
where the integration domain of x and y is the whole AdS boundary $\partial\mathcal{M}=A\bar{A}$. Here we have considered all the PEE threads and their weights instead of only those connecting $A$ and $\bar{A}$. The factor $1/2$ appears as we count both $\mathcal{I}(\textbf{x},\textbf{y})$ and $\mathcal{I}(\textbf{y},\textbf{x})$ in the integration, hence every PEE thread has been counted twice. Conducting the integration (22) is formidable since it is impossible to trace the weight $\omega_{\Sigma_{A}}(\textbf{x},\textbf{y})$ of all the PEE threads. The formula (22) will be quite useful when we relate our results to the Crofton formula.